A flywheel turns at 456 rpm. Compute the angular speed (in rad/s) at any point in the wheel.


1
Expert's answer
2021-05-26T10:33:37-0400
"\omega=456\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot\dfrac{2\pi\ rad}{1\ rev}=47.7\ \dfrac{rad}{s}."